# Speech emotion recognition from voice messages recorded in the wild

### v1

@@ VI Discussion
There is limited research on evaluating model fairness, particularly in the context of pre-trained models [27]. Our focus here is on gender, given the insufficient information for other attributes considered in the reference databases. We measure fairness by calculating the difference between the UA for the male speakers and the UA for the female speakers (see Fig. 7). For the EMOVOME database, models trained using expert labels exhibit a notable bias towards males, as the UA is around 10% higher for males in valence and arousal prediction and 1.7% in emotion categories prediction. The use of non-expert labels resulted in an increase in UA for males of 4.6% valence, but it was 1.9% higher for females on arousal. Interestingly, the combined label yielded the most similar results for both genders, with 0.3% for valence and 1.4% for arousal. In the case of IEMOCAP, again, UA was higher for male speakers in valence (4%) but lower in arousal (-3.4%). For categories, the UA for males was +5.9% compared to females. Finally, RAVDESS presents the highest difference between both genders, with the UA for females being 29.7% higher for females compared to males. For arousal and valence, the results are also higher for females (1% and 6.5%, respectively).
@@
Overall, SER models obtain better test results for male speakers in EMOVOME, following previous studies in Spanish databases [21, 42]. In the reference databases, IEMOCAP aligns with the observed trend, while RAVDESS shows the opposite results. However, it’s important to note that both databases have a limited number of speakers in the test set (two for IEMOCAP and five for RAVDESS), so no significant conclusions can be drawn from these results.
VI-E Limitations and future work
@@
This section addresses the identified limitations, offering insights into areas for potential improvement in future research. Firstly, enhancing the annotation process of EMOVOME samples in emotion categories may involve introducing new raters to alleviate potential individual biases that could impact SER models. Similarly, expanding the pool of non-expert labels with more raters may help mitigate bias and achieve a more balanced data distribution regarding arousal. Additionally, other databases, such as EmoSpanishDB or MOUD, could be explored to increase the number of training samples and assess potential improvements in model performance. Furthermore, in the creation of SER models based on pre-trained models, the current use of average time pooling as an aggregation approach may exhibit suboptimal performance, particularly noticeable in the EMOVOME database, where audio duration exhibits significant variability. Future research endeavors will focus on refining time aggregation methods and exploring alternative techniques to address these challenges effectively.
VII Conclusions and future work
@@
A comprehensive study was conducted to assess the influence of different properties of EMOVOME, a natural speech database representative of real-world settings created for this work, on the performance of speaker-independent SER models. Superior results were achieved with state-of-the-art pre-trained transformer-based models compared to baseline models based on acoustic features. However, these results demonstrated lower performance compared to the SER models trained on the acted RAVDESS database. For the elicited IEMOCAP database, the prediction of emotion categories outperformed EMOVOME, but similar results were obtained for predicting valence and arousal. Notably, we found variations in EMOVOME results depending on the labels provided by different annotators, with superior outcomes observed when utilizing combined labels from both expert and non-experts. Interestingly, this combined label also yielded the most equitable results when assessing gender fairness, even though SER models generally performed better for male speakers.
@@ Acknowledgments
This work was supported by the European Union’s Horizon 2020 funded project “HELIOS: A Context-aware Distributed Social Networking Framework” (No 825585), by the Universitat Politècnica de València (PAID-10-20), by the Generalitat Valenciana (ACIF/2021/187 and PROMETEO/2020/024), by the Spanish Government (BEWORD PID2021-126061OB-C41) and by the Spanish Ministry of Science and Innovation for the DIPSY project (TED2021-131401B-C21).

### v3

@@ VII-C Fairness evaluation
There is limited research on evaluating model fairness, particularly in the context of pre-trained models [46]. Our focus here is on gender, given the insufficient information for other attributes considered in the reference datasets (see Fig. 7). For EMOVOME dataset, models trained using expert labels exhibit a notable bias towards males, as the UA is around 10% higher for males in valence and arousal prediction and 1.7% in emotion categories prediction. The use of non-expert labels resulted in an increase in UA for males of 4.6% valence, but it was 1.9% higher for females on arousal. Interestingly, the combined label yielded the most similar results for both genders, with 0.3% for valence and 1.4% for arousal. In the case of IEMOCAP, again, UA was higher for male speakers in valence (4%) but lower in arousal (-3.4%). For categories, the UA for males was +5.9% compared to females. Finally, RAVDESS presents the highest difference between both genders, with the UA for females being 29.7% higher for females compared to males. For arousal and valence, the results are also higher for females (1% and 6.5%, respectively).
@@
Overall, SER models obtain better test results for male speakers in EMOVOME, following previous studies in Spanish datasets [30, 38]. In the reference datasets, IEMOCAP aligns with the observed trend, while RAVDESS shows the opposite results. However, it’s important to note that both datasets have a limited number of speakers in the test set (two for IEMOCAP and five for RAVDESS), so no significant conclusions can be drawn from these results.
VII-D Limitations and future work
@@
Firstly, while we manually screened for ad hoc recordings, complete control over this factor remains challenging, necessitating reliance on participant integrity. Secondly, improving the annotation process of EMOVOME samples within emotion categories by introducing new raters could alleviate potential individual biases that impact SER models. Similarly, expanding the pool of non-expert labels with more raters may help mitigate bias and achieve a more balanced data distribution regarding arousal. Additionally, other datasets, such as EmoSpanishDB or MOUD, could be explored to increase the number of training samples and assess potential improvements in model performance. Furthermore, in the creation of SER models based on pre-trained models, the current use of average time pooling as an aggregation approach may exhibit suboptimal performance, particularly noticeable in the EMOVOME dataset, where audio duration exhibits significant variability. Future research endeavors will focus on refining time aggregation methods and exploring alternative techniques to address these challenges effectively.
VIII Conclusions
@@
We developed and publicly released EMOVOME, the first public dataset with realistic emotions from spontaneous conversations in a real-life scenario. We developed speaker-independent SER models using EMOVOME, and compared the results with IEMOCAP and RAVDESS. Superior results were achieved with pre-trained transformer-based models compared to baseline models based on acoustic features. However, EMOVOME results demonstrated lower performance compared to the acted RAVDESS dataset. For the elicited IEMOCAP dataset, the prediction of emotion categories outperformed EMOVOME, but similar results were obtained for predicting valence and arousal. A comprehensive study was also conducted to assess the influence of different properties of EMOVOME on SER performance. Notably, we found variations depending on the labels provided by different annotators, with superior outcomes observed when utilizing combined labels from both expert and non-experts. Interestingly, this combined label also yielded the most equitable results when assessing gender fairness. The results highlight the importance of natural datasets for predicting genuine emotions, emphasizing the need for speech from real-life scenarios to accurately assess the challenges associated with authentic environments.
@@ Acknowledgments
This work was supported by the European Union’s Horizon 2020 funded project HELIOS (ID 825585), by the Universitat Politècnica de València (PAID-10-20), by the Generalitat Valenciana (ACIF/2021/187), and by the Spanish Ministry of Science and Innovation for the BEWORD project (PID2021-126061OB-C41) and for the DIPSY project (TED2021-131401B-C21).